X~N(0,1) , find (i) x s. t. Pr (-x<X<x)=0.95
(ii) x s.t. Pr. (-x<X<x) = 0.90.

I le <x)=0.95. X2N(0,1), find (i) x sit. Pr(-x < X (ii) x s.t. Pr(-x < X <x)=0.90. Ti l a botter in a laptop con
Answer #1

(i) P(-x<X<x)=0.95 First we need to find the z values for the given probability P(-z<Z<z)=0.95 P(Z<z)-P(Z<-z)=0.95 P(Z<z)-[1-(ii) P(-x<X<<x)=0.90 First we need to find the z values for the given probability P(-z<Z<z)=0.90 P(Z<z)-P(Z<-z)=0.90 P(Z<z)-[